A bicyclist and a hiker leave the same place at the same time and travel in the same direction. The bicyclist travels three times as fast as the hiker. At the end of 3hr they are 24 miles apart. How fast does the hiker travel?
Answer by stanbon(75887)

A bicyclist and a hiker leave the same place at the same time and travel in the same direction. The bicyclist travels three times as fast as the hiker. At the end of 3hr they are 24 miles apart. How fast does the hiker travel?
-------------
Bike DATA: rate = 3x mph ; time = 3 hrs ; distance r*t = 9x miles
-----
Hike DATA: rate = x mph ; time = 3 hrs ; distance = r*t = 3x miles
------
Equation:
distance - distance = 24 miles
9x - 3x = 24
6x = 24
x = 4 mph (hiker rate)
3x = 12 mph (biker rate)
============================
Cheers,
Stan H.
===============